OpenStudy (anonymous):

what is the minimum value of 7^(x^2-4x+7)

OpenStudy (saifoo.khan):

\[\Large 7^{x^2-4x+7}\]Like this?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

@mathmate @UnkleRhaukus

OpenStudy (saifoo.khan):

Minimum value should be there where dy/dx = 0.

OpenStudy (anonymous):

what do you mean by the "d"?

OpenStudy (saifoo.khan):

In other words minimum value occurs where the derivative of the function is equal to zero.

OpenStudy (anonymous):

isn't derivatives calculus? i haven't learned that yet and i don't think we're expected to know it to solve this problem

OpenStudy (anonymous):

haha i was thinking about it and wouldn't it be one over infinity? because if x^2-4x+7 was negative infinity than 7^(x^2-4x+7) would be 1/infinity

OpenStudy (mathmate):

Hint: 7^x is an increasing function. So if we minimize x, 7^x is also a minimum.

OpenStudy (mathmate):

x^2-4x+7 cannot be at -inf. It is a parabola concave upwards.!

OpenStudy (saifoo.khan):

And are we going to try that by trial and error? @mathmate

OpenStudy (kinggeorge):

No, we need to simply find the vertex of the parabola to find the minimum value of the parabola.

OpenStudy (saifoo.khan):

Howw??

OpenStudy (mathmate):

Good point!] @livelaughlilz can you take it from here?

OpenStudy (mathmate):

Completing the squares will do the job.

OpenStudy (anonymous):

yeah the vertex of the parabola is (2,3)

OpenStudy (anonymous):

(x-2)^2=y-3

OpenStudy (mathmate):

x^2-4x+7=(x-2)^2+3 so the vertex is at (2,3)

OpenStudy (mathmate):

Good job livelaughlilz! Speed+accuracy!

OpenStudy (mathmate):

I like his movies! :)

OpenStudy (anonymous):

oh so the answer is 343. the lowest value of x is 2, and if you plug it in you get 7^3=343

OpenStudy (saifoo.khan):

Yes^ @livelaughlilz

OpenStudy (mathmate):

7^2 = ???

OpenStudy (saifoo.khan):

@mathmate : haha, nice. My brother just love his movies. i don't. :/

OpenStudy (mathmate):

Nevermind, I was out of my mind. Yes, 7^3=343

OpenStudy (saifoo.khan):

You were thinking about Jackie Chan. :D

OpenStudy (kinggeorge):

As a sidenote @mathmate, you can find the x-coordinate of the vertex using the formula \[v_x=\frac{-b}{2a}\]

OpenStudy (mathmate):

Thank you for the defence, even though it was a lame one! Good thinking! :)

OpenStudy (saifoo.khan):

I simply love this method. Satellite73 taught me this. -b/2a

OpenStudy (anonymous):

to add onto @KingGeorge the y-coordinate of the vertex is -[(b^2-4ac)/(4a)]

OpenStudy (mathmate):

Thank you @KingGeorge, my memory is limited and sometimes defective. But thanks for the shortcut anyway!

OpenStudy (anonymous):

you can figure all that out by completing the square of y=ax^2+bx+c

OpenStudy (anonymous):

and making y=ax^2+bx+c into vertex form

OpenStudy (mathmate):

Actually I know the -b/2a part, that's how I get to do the completing the square. It's just I never managed to memorize the -[(b^2-4ac)/(4a)] part.

OpenStudy (anonymous):

well b^2-4ac is also the discriminant...

OpenStudy (saifoo.khan):

@livelaughlilz is a boss at this.

OpenStudy (mathmate):

Perhaps from this time on I will remember. Practice makes perfect! :)

OpenStudy (mathmate):

Actually what I do is finding it by c-(b/2a)^2, which is exactly what we do when we complete the square.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i'm going to close this question now. thanks everyone!

OpenStudy (mathmate):

See you all!